# Constraining electromagnetic couplings of ultralight scalars from compact stars

### v1

@@ VII Conclusions and Discussions
The interaction of the scalar with the EM fields modifies Maxwell’s EM equations and gives rise to scalar-induced electric and magnetic fields. The dispersion relation of the EM radiation (photon) emitted by the star is also modified during its propagation through the long-range scalar field. This would result in a wavelength-dependent apparent redshift of photons emitted by the star. The measurement of this wavelength dependence, combined with the knowledge of the redshift of host galaxy, can lead to the determination of $g_{\phi\gamma\gamma}$. With the currently possible precision on redshifts ($\delta z\sim 10^{-4}$), one can be sensitive to $g_{\phi\gamma\gamma}\sim 10^{-16}$ GeV-1. With future precision atomic clocks that may be able to measure a specific spectral line from the magnetar with a precision of $\delta z=\Delta k/k\sim 10^{-18}$, the sensitivity to $g_{\phi\gamma\gamma}\sim 10^{-23}$ GeV-1 may be obtained.
@@
In a binary pulsar system where both of the compact stars give rise to scalar fields, the stars experience a long-range scalar-mediated force in addition to gravity, arising from the interaction between the scalar fields and EM fields sourced by the two stars. The scalar-mediated force may be mimicked by a change in the product of masses of the two stars. However, if independent information about the two stars and the distance between them is available – for example, from their interaction with a third body gravitationally bound to them but without a large EM field – then it would be possible to detect this force or to constrain its value. Using the parameters of the pulsar binary system PSR J0737-3039, if the masses of the two stars and the distance between them is known to a precision of $0.05\%$, we find that a constraint of $g_{\phi\gamma\gamma}\lesssim 8\times 10^{-6}~{}\mathrm{GeV}^{-1}$ on the scalar-photon coupling may be obtained. This constraint would be valid for a scalar mass of $m_{\phi}\lesssim 2.2\times 10^{-16}~{}\mathrm{eV}$ so that the range of the force is more than the distance between the two stars in this binary system. This bound is weaker than that from the fifth force measurements by several orders of magnitude. However, it is inversely proportional to the square of the magnetic field at the surface of each star, and future measurements of a binary magnetar system with a high magnetic field $(B_{0}\sim 10^{17}~{}\mathrm{G})$ could improve it to $g_{\phi\gamma\gamma}\lesssim 10^{-18}~{}\mathrm{GeV}^{-1}$ if the masses of the stars and the distance between them are known to $\sim 0.05\%$.
@@
If the background EM fields are time-dependent, the scalar field will also be time-dependent and hence will radiate. This scalar radiation will lead to an additional spin-down of the star. The spin-down luminosity, or the rate of change of rotational energy of the star, is a measurable quantity and can be obtained from the measurements of the spin period and its derivative. The scalar spin-down luminosity increases with increasing surface magnetic field, radius, and spin frequency of the star. We analyze the data on the Crab pulsar, SGR 1806-20, and GRB 080905A, and obtain the strongest bound on the scalar-photon coupling from the measurement of the spin-down luminosity of GRB 080905A as $g_{\phi\gamma\gamma}\lesssim 2\times 10^{-16}~{}\mathrm{GeV}^{-1}$ for $m_{\phi}\lesssim 4.2\times 10^{-13}~{}\mathrm{eV}$.
@@
The constraints discussed here from various observations can be further improved with enhanced sensitivity of detection and by focusing on stars with high surface magnetic fields, larger radii, and higher spin frequencies. The strongest constraint on the scalar-photon coupling comes from the measurements of the rate of change of the spin period due to the EM radiation. This constraint is stronger than the astrophysical bound from the globular clusters. For extremely low values of $m_{\phi}$, i.e., for $m_{\phi}\lesssim 1.5\times 10^{-22}~{}\mathrm{eV}$, it is also stronger than the bound set by the fifth-force experiments.
The reason why our method yields stronger bounds than that from the fifth-force measurements at ultralight masses is as follows. The fifth-force experiments investigate the derivative of a generic Yukawa potential across small and large length scales to identify any deviations from standard gravity. In the very long-range limit, the mass of the mediator approaches zero, making the Yukawa potential indistinguishable from the standard Newtonian potential. In our method, the constraints on the coupling are valid when the scalar mass is smaller than the relevant inverse-distance scale in the observed system: the inverse of the distance between the binaries, the inverse radius of the compact star, or the spin frequency of the compact star. Therefore, our results stay valid as $m_{\phi}$ goes to zero.
Ultralight scalar particles are predicted in many theories, and investigating their possible existence and properties is important, irrespective of whether they form a significant amount of DM in the Universe. These particles can give rise to interesting observable effects in astrophysics as well as in laboratory tests of the equivalence principle. The bounds obtained by us in this paper (see Fig. 1), using the available data on pulsars and magnetars, are not yet competitive with precision Eöt-Wash and MICROSCOPE experiments. However, future observations of astrophysical systems by detectors with higher sensitivities, and with precision clocks, can improve these bounds by orders of magnitude.
@@ Acknowledgements
The authors would like to thank Shadab Alam, Girish Kulkarni, Jamie Mcdonald and Nicholas Rodd for useful discussions. T. K. P. would also like to thank the Galileo Galilei Institute for Theoretical Physics for the hospitality and the INFN for partial support during the completion of this work, and COST Actions COSMIC WISPers CA21106 and BridgeQG CA23130, supported by COST (European Cooperation in Science and Technology). The work of A.D. is supported by the Department of Atomic Energy, Government of India, under Project Identification Number RTI 4002.
Appendix A Long-range scalar field outside a compact star in Kerr background
@@
$\mathbf{B}^{\mathrm{in}}_{(r=R)}=B_{0}\Big{(}\cos\theta\hat{r}+\frac{\sin\theta}{2}\hat{\theta}\Big{)}.$ (33)
@@
$\mathbf{E}^{\mathrm{in}}_{(r=R)}=B_{0}\Omega R\sin\theta\Big{(}\frac{\sin\theta}{2}\hat{r}-\cos\theta\,\hat{\theta}\Big{)}.$ (35)
@@
Since the star is rotating, the geometry outside of the star is described by the Kerr metric, given as
$ds^{2}=\frac{\rho^{2}}{\Delta}dr^{2}+\rho^{2}d\theta^{2}+\Big{[}(r^{2}+a^{2})\sin^{2}\theta+\frac{2Mr}{\rho^{2}}a^{2}\sin^{4}\theta\Big{]}d\varphi^{2}-\frac{4Mr}{\rho^{2}}a\sin^{2}\theta d\varphi dt-\Big{(}1-\frac{2Mr}{\rho^{2}}\Big{)}dt^{2},$ (38)
in Boyer-Lindsquist coordinates. Here, $\Delta=r^{2}-2Mr+a^{2}$, $\rho^{2}=r^{2}+a^{2}\cos^{2}\theta$ and $a=J/M$, where $J$ is the spin angular momentum of the star and $M$ is the stellar mass. To solve Eq. 5 in the Kerr background when $\Omega$ is not very small, it is essential to account for terms proportional to $\mathcal{O}(a^{2})$ for large $\Omega$. This is because the term $(\mathbf{B}^{2}-\mathbf{E}^{2})$ in Eq. 3 contains $\Omega^{2}$, making it impossible to neglect higher-order contributions in $a^{2}$. Hence, the equation of motion of the scalar field in the Kerr background can be written as
$\begin{split}\frac{1}{(r^{2}+a^{2}\cos^{2}\theta)}\frac{\partial}{\partial r}\Big{[}(r^{2}-2Mr+a^{2})\frac{\partial}{\partial r}\phi(r,\theta)\Big{]}+\frac{1}{(r^{2}+a^{2}\cos^{2}\theta)\sin\theta}\frac{\partial}{\partial\theta}\Big{[}\sin\theta\frac{\partial}{\partial\theta}\phi(r,\theta)\Big{]}\\ =-g_{\phi\gamma\gamma}\frac{B^{2}_{0}R^{6}}{4r^{6}}(3\cos^{2}\theta+1)+g_{\phi\gamma\gamma}\frac{B^{2}_{0}\Omega^{2}R^{10}}{4r^{8}}(5\cos^{4}\theta-2\cos^{2}\theta+1).\end{split}$ (39)
To solve the above inhomogeneous differential equation (Eq. 39), we use the Green’s function method. The source term is given as
@@
We can write Eq. 41 in the Kerr background, for a point source at $r=b$ and $\theta_{0}=\varphi_{0}=0$, as
$\begin{split}\frac{1}{(r^{2}+a^{2}\cos^{2}\theta)}\frac{\partial}{\partial r}\Big{[}(r^{2}-2Mr+a^{2})\frac{\partial G}{\partial r}\Big{]}+\frac{1}{(r^{2}+a^{2}\cos^{2}\theta)\sin\theta}\frac{\partial}{\partial\theta}\Big{[}\sin\theta\frac{\partial G}{\partial\theta}\Big{]}\\ =-\frac{\delta(r-b)\delta(\cos\theta_{0}-1)\delta(\varphi_{0})}{(r^{2}+a^{2}\cos^{2}\theta)}.\end{split}$ (43)
@@
$\frac{\partial}{\partial r}\Big{[}(r^{2}-2Mr+a^{2})\frac{\partial R_{l}}{\partial r}\Big{]}-l(l+1)R_{l}=0.$ (45)
Therefore, considering that the scalar field is finite at $r\rightarrow\infty$ and at $r\rightarrow M+\sqrt{M^{2}-a^{2}}$, and continuous at $r=b$, we obtain the solution of the Green’s function as Campbell:1991kz
$G(r,\theta)$ $=$ $\sum^{\infty}_{l=0}C_{l}P_{l}\Big{(}\frac{b-M}{\sqrt{M^{2}-a^{2}}}\Big{)}Q_{l}\Big{(}\frac{r-M}{\sqrt{M^{2}-a^{2}}}\Big{)}P_{l}(\cos\theta),~{}~{}~{}r>b\ ,$ (46)
$=$ $\sum^{\infty}_{l=0}C_{l}Q_{l}\Big{(}\frac{b-M}{\sqrt{M^{2}-a^{2}}}\Big{)}P_{l}\Big{(}\frac{r-M}{\sqrt{M^{2}-a^{2}}}\Big{)}P_{l}(\cos\theta),~{}~{}~{}r<b\ ,$
where $P_{l}$ and $Q_{l}$ denote the Legendre polynomials of degree $l$ of first and second kind, respectively, and $C_{l}=(2l+1)/(4\pi\sqrt{M^{2}-a^{2}})$.
@@
where $r_{s}=M+\sqrt{M^{2}-a^{2}}$. Since the source term does not depend on $\varphi_{0}$, we can immediately perform the integration for $\varphi_{0}$ and write Eq. 47 as
[…]
$\phi(r)\approx-\frac{g_{\phi\gamma\gamma}B^{2}_{0}\Omega^{2}R^{10}}{480M^{5}r}+\frac{g_{\phi\gamma\gamma}B_{0}^{2}R^{6}}{576M^{4}r}\Big{(}\frac{15a^{2}}{M}+12M\Big{)}+\frac{7g_{\phi\gamma\gamma}B^{2}_{0}R^{6}a^{2}}{4800M^{5}r}+\mathcal{O}\Big{(}\frac{1}{r^{2}}\Big{)},$ (49)
where $\Omega=a/2M(M+\sqrt{M^{2}-a^{2}})$. The dominant term of the scalar field is the monopole term $(l=0)$ and we can write the scalar field configuration as $\phi(r)\approx Q_{\phi}^{\mathrm{K}}/r$, where $Q_{\phi}^{\mathrm{K}}$ is the scalar charge, defined as
$Q_{\phi}^{\mathrm{K}}=-\frac{g_{\phi\gamma\gamma}B^{2}_{0}\Omega^{2}R^{10}}{480M^{5}}+\frac{g_{\phi\gamma\gamma}B^{2}_{0}R^{6}}{576M^{4}}\Big{(}\frac{15a^{2}}{M}+12M\Big{)}+\frac{7g_{\phi\gamma\gamma}B^{2}_{0}R^{6}a^{2}}{4800M^{5}}.$ (50)
In the limit $a\rightarrow 0$ and $\Omega R\ll 1$, Eqs. 49 and 50 reduce to Eqs. 8 and 9 respectively.
The scalar-induced magnetic field can be obtained by solving Eq. 13 in the Kerr background as
$\begin{split}\mathbf{B}_{\phi}(r,\theta)\approx-\frac{g_{\phi\gamma\gamma}B_{0}R^{3}}{4M^{2}}\Big{(}\frac{Q_{\phi}^{\mathrm{K}}}{3}+\frac{59Q^{\mathrm{eff}}_{\phi}a^{2}}{400M^{2}}\Big{)}\frac{\cos\theta}{r^{2}}\hat{r}-\frac{g_{\phi\gamma\gamma}B_{0}R^{3}\pi}{64M^{3}r}\Big{(}Q_{\phi}^{\mathrm{K}}+\frac{103Q^{\mathrm{eff}}_{\phi}a^{2}}{80M^{2}}\Big{)}\hat{\theta},\end{split}$ (51)
where $Q_{\phi}^{\mathrm{eff}}$ is given in Eq. 9. The limiting scenario where $\Omega R\ll 1$ from Eq. 51 is given by Eq. 14.

### v2

@@ VII Conclusions and Discussions
We work in the context of a theory with an effective scalar-photon coupling $g_{\phi\gamma\gamma}$. Such a coupling would typically arise from a UV theory with a mass scale $\Lambda$. In our work, $g_{\phi\gamma\gamma}$, and hence this $\Lambda$, are effective parameters and we are agnostic about their origin. The bounds obtained on $g_{\phi\gamma\gamma}$ can be translated to the bounds on $\Lambda$ in the context of a UV-complete theory, but it is beyond the scope of our work. The mass of the ultralight scalar, $m_{\phi}$, may also be related to $\Lambda$ in the context of a UV-complete theory. However, in the phenomenological work of ours, it is simply a free mass parameter.
@@
The interaction of the scalar with the EM fields modifies Maxwell’s EM equations and gives rise to scalar-induced electric and magnetic fields. The dispersion relation of the EM radiation (photon) emitted by the star is also modified during its propagation through the long-range scalar field. This would result in a wavelength-dependent apparent redshift of photons emitted by the star. The measurement of this wavelength dependence, combined with the knowledge of the redshift of host galaxy, can lead to the determination of $g_{\phi\gamma\gamma}$. With the currently possible precision on redshifts ($\delta z\sim 10^{-4}$), one can be sensitive to $g_{\phi\gamma\gamma}\sim 10^{-15}$ GeV-1, using the benchmark GRB 080905A. With future precision atomic clocks that may be able to measure a specific spectral line from the magnetar with a precision of $\delta z\sim\Delta k/k\sim 10^{-18}$, the sensitivity to $g_{\phi\gamma\gamma}\sim 10^{-19}~\mathrm{GeV}^{-1}$ may be obtained. With precision measurements of low-frequency photon wavenumbers using entangled quantum clock networks, the sensitivity to the scalar-photon coupling $g_{\phi\gamma\gamma}$ would be competitive with the existing bounds.
@@
In a binary pulsar system where both of the compact stars give rise to scalar fields, the stars experience a long-range scalar-mediated force in addition to gravity, arising from the interaction between the scalar fields and EM fields sourced by the two stars. The scalar-mediated force may be mimicked by a change in the masses of the two stars. However, if independent information about the masses of the two stars and the distance between them is available – for example, from their interaction with a third body gravitationally bound to them but without a large EM field – then it would be possible to detect this force or to constrain its value. Using the parameters of the pulsar binary system PSR J0737-3039, if the masses of the two stars and the distance between them is known to a precision of $0.05\%$, we find that a constraint of $g_{\phi\gamma\gamma}\lesssim 8\times 10^{-6}~\mathrm{GeV}^{-1}$ on the scalar-photon coupling may be obtained. This constraint would be valid for a scalar mass of $m_{\phi}\lesssim 2.2\times 10^{-16}~\mathrm{eV}$ so that the range of the force is more than the distance between the two stars in this binary system. This bound is weaker than that from the fifth force measurements by several orders of magnitude. However, it is inversely proportional to the square of the magnetic field at the surface of each star, and future measurements of a binary magnetar system with a high magnetic field $(B_{0}\sim 10^{17}~\mathrm{G})$ could improve it to $g_{\phi\gamma\gamma}\lesssim 10^{-18}~\mathrm{GeV}^{-1}$ if the masses of the stars and the distance between them are known to $\sim 0.05\%$.
@@
If the background EM fields are time-dependent, the scalar field will also be time-dependent and hence will radiate. This scalar radiation will lead to an additional spin-down of the star. The spin-down luminosity, or the rate of change of rotational energy of the star, is a measurable quantity and can be obtained from the measurements of the spin period and its derivative. The scalar spin-down luminosity increases with increasing surface magnetic field, radius, and spin frequency of the star. We analyze the data on the Crab pulsar, SGR 1806-20, and GRB 080905A, and obtain the strongest bound on the scalar-photon coupling from the measurement of the spin-down luminosity of GRB 080905A as $g_{\phi\gamma\gamma}\lesssim 2\times 10^{-14}~\mathrm{GeV}^{-1}$ for $m_{\phi}\lesssim 4.2\times 10^{-13}~\mathrm{eV}$.
@@
The constraints discussed here from various observations can be further improved with enhanced sensitivity of detection and by focusing on stars with high surface magnetic fields, larger radii, and higher spin frequencies. The strongest constraint on the scalar-photon coupling comes from the measurements of the rate of change of the spin period due to the EM radiation. Note that the constraints presented on $g_{\phi\gamma\gamma}$ are actually on its magnitude. The expressions for the scalar-induced apparent redshift of photons, the scalar-induced magnetic field, the energy loss rate from pulsars due to scalar radiation, and the scalar-mediated fifth-force between two pulsars all depend on even powers of the coupling $g_{\phi\gamma\gamma}$. As a result, the constraints derived on $g_{\phi\gamma\gamma}$ are insensitive to the sign of the coupling.
Our analysis of electromagnetic radiation due to scalar-induced magnetic field from GRB080905A yields the most stringent astrophysical constraint on $g_{\phi\gamma\gamma}$ to date, improving the previous astrophysical bound from globular clusters by eight orders of magnitude. At extremely low masses, i.e. for $m_{\phi}\lesssim 10^{-22}~\mathrm{eV}$, our bound are stronger than that from the fifth-force experiments. The existing bounds from the atomic clock experiments assume the scalar to be the DM, and hence do not directly apply to our scenario. The MICROSCOPE bound, obtained through the measurement of the variation of nuclear binding energy induced by the change in the fine structure constant, is stronger. However, our bounds are obtained from astrophysical observations, without involving detailed nuclear physics considerations, and are therefore complementary to this bound. In future, developments such as highly sensitive nuclear or space-based entangled clock systems distributed over large baselines, combined with the detection of low frequency photon signals by instruments like LOFAR and SKA, could enable precise measurements of deviations in the photon wavenumber. This would potentially yield limits that are competitive with those from current equivalence principle experiments.
Ultralight scalar particles arise in many theoretical models and motivate diverse experimental searches, regardless of their contribution to DM. We have explored their observable implications within a minimal framework, characterized solely by the scalar mass $m_{\phi}$ and its coupling to photons $g_{\phi\gamma\gamma}$, without invoking hidden-sector assumptions. We point out for the first time that such scalars generate a long-range monopole field outside magnetized stars, and show that, unlike axions, they do not induce birefringence. Future precision measurements of pulsars and magnetars could pave the way for a better understanding of the physics of ultralight scalars, in addition to the astrophysics of these compact objects.
@@ Acknowledgements
The authors would like to thank Shadab Alam, Sumanta Chakraborty, Girish Kulkarni, Jamie Mcdonald, Arunava Mukherjee and Nicholas Rodd for useful discussions. The authors would also like to thank the anonymous referee for their fruitful comments and suggestions. T. K. P. would also like to thank the Galileo Galilei Institute for Theoretical Physics and ICTP for the hospitality and the INFN for partial support during the completion and finalization of this work, and COST Actions COSMIC WISPers CA21106 and BridgeQG CA23130, supported by COST (European Cooperation in Science and Technology). The work of A.D. is supported by the Department of Atomic Energy, Government of India, under Project Identification Number RTI 4002.
Appendix A Long-range scalar field outside a compact star
@@
[…]
@@
[…]
@@
To solve Eq. 5 when $\Omega$ is not negligibly small, it is crucial to include the contribution from $\mathbf{E}^{2}$, since $\mathbf{E}^{2}$ in Eq. 3 contains terms proportional to $\Omega^{2}$. Therefore, we use the Green’s function method in solving the inhomogeneous differential equation Eq. 5. The source term is given as
@@
We can write Eq. 45 for a point source in a Schwarzschild background at $r=b$ and $\theta_{0}=\varphi_{0}=0$, as
$\begin{split}\frac{1}{r^{2}}\frac{\partial}{\partial r}\Big[(r^{2}-2Mr)\frac{\partial G}{\partial r}\Big]+\frac{1}{r^{2}\sin\theta}\frac{\partial}{\partial\theta}\Big[\sin\theta\frac{\partial G}{\partial\theta}\Big]=-\frac{\delta(r-b)\delta(\cos\theta_{0}-1)\delta(\varphi_{0})}{r^{2}}.\end{split}$ (47)
@@
$\frac{\partial}{\partial r}\Big[(r^{2}-2Mr)\frac{\partial R_{l}}{\partial r}\Big]-l(l+1)R_{l}=0.$ (49)
Therefore, considering that the scalar field is finite at $r\rightarrow\infty$ and at $r\rightarrow 2M$, and continuous at $r=b$, we obtain the solution of the Green’s function as Campbell:1991kz
$G(r,\theta)$ $=$ $\sum^{\infty}_{l=0}C_{l}P_{l}\Big(\frac{b-M}{M}\Big)Q_{l}\Big(\frac{r-M}{M}\Big)P_{l}(\cos\theta),~~~r>b\ ,$ (50)
$=$ $\sum^{\infty}_{l=0}C_{l}Q_{l}\Big(\frac{b-M}{M}\Big)P_{l}\Big(\frac{r-M}{M}\Big)P_{l}(\cos\theta),~~~r<b\ ,$
where $P_{l}$ and $Q_{l}$ denote the Legendre polynomials of degree $l$ of first and second kind, respectively, and $C_{l}=(2l+1)/(4\pi M)$.
@@
$\phi(r,\theta)=-\int^{\infty}_{r_{s}}db\int^{\pi}_{0}d\theta_{0}\int^{2\pi}_{0}d\varphi_{0}b^{2}\sin\theta_{0}G(r,\theta,\varphi,b,\theta_{0},\varphi_{0})J(b,\theta_{0},\varphi_{0}),$ (51)
where $r_{s}=2M$. Since the source term does not depend on $\varphi_{0}$, we can immediately perform the integration for $\varphi_{0}$ and write Eq. 51 as
$\begin{split}\phi(r,\theta)=-\sum_{l=0}^{\infty}\frac{2l+1}{2M}\int^{r}_{2M}db\int^{\pi}_{0}d\theta_{0}b^{2}\sin\theta_{0}P_{l}\Big(\frac{b-M}{M}\Big)Q_{l}\Big(\frac{r-M}{M}\Big)P_{l}(\cos\theta)P_{l}(\cos\theta_{0})J(b,\theta_{0})\\ -\sum^{\infty}_{l=0}\frac{2l+1}{2M}\int^{\infty}_{r}db\int^{\pi}_{0}d\theta_{0}b^{2}\sin\theta_{0}Q_{l}\Big(\frac{b-M}{M}\Big)P_{l}\Big(\frac{r-M}{M}\Big)P_{l}(\cos\theta)P_{l}(\cos\theta_{0})J(b,\theta_{0}).\end{split}$ (52)
$\phi(r)\approx-\frac{g_{\phi\gamma\gamma}B^{2}_{0}\Omega^{2}R^{10}}{480M^{5}r}+\frac{g_{\phi\gamma\gamma}B_{0}^{2}R^{6}}{48M^{3}r}+\mathcal{O}\Big(\frac{1}{r^{2}}\Big).$ (53)
The dominant term of the scalar field is the monopole term $(l=0)$ and we can write the scalar field configuration as $\phi(r)\approx Q_{\phi}^{\mathrm{K}}/r$, where $Q_{\phi}^{\mathrm{K}}$ is the scalar charge, defined as
$Q_{\phi}^{\mathrm{K}}=-\frac{g_{\phi\gamma\gamma}B^{2}_{0}\Omega^{2}R^{10}}{480M^{5}}+\frac{g_{\phi\gamma\gamma}B^{2}_{0}R^{6}}{48M^{3}}.$ (54)
In the limit $\Omega R\ll 1$, Eqs. 53 and 54 reduce to Eqs. 7 and 8 respectively.
The scalar-induced magnetic field in the large $\Omega$ limit can similarly be obtained by solving Eq. 13 as
@@
where $Q_{\phi}^{\mathrm{K}}$ is given in Eq. 54. The limiting scenario where $\Omega R\ll 1$ from Eq. 55 is given by Eq. 14. Therefore, the scalar-induced magnetic field is dipolar $(l=1)$ along the radial direction and monopolar $(l=0)$ along the angular direction.